Two objects A and B are separated by a distance 100 m. They decide to move towards another point C in the same direction simultaneously with velocities 20 m/s and 40m/s with A ahead of B. What is the distance between them 2 s after they start?


1
Expert's answer
2021-08-02T08:42:00-0400
"s=D-(v_1-v_2)t\\s=100-(-20+40)2=60\ m"